The demand function Q(P) and cost functions C(Q) of a commodity are given by the equations:

Q=12000−60P

C(Q)=10000+4Q,

where P and Q are the price and quantity, respectively. 

The total revenue function TR in terms of P is


a. TR=12 000−60P.

b. TR=P(12 000−60P^2).

c.  TR=12 000P−60P^2.

d.  TR=12 000+60P^2.



1
Expert's answer
2022-04-25T17:22:11-0400

The total revenue function (TR), in terms of P

TR = Q * P = 12000P - 60P2; Q =12 000 - 60 P

answer:c